can you show me how to do
5x-1 over 2 = 4 - x over 2

Answer by stanbon(75887)

5x-1 over 2 = 4 - x over 2
-----------------------------
(5x-1)/2 = 4 - (x/2)
Multiply thru by 2 to get:
5x-1 = 8 - x
Combine:
6x = 9
x = 3/2
=============
Cheers,
Stan H.